The line (x-y)/3=(y+1)/4=(z-2)/3 passes through the point (4,2,3). Is it true or false. Give reasons for your answer


1
Expert's answer
2021-02-24T07:07:38-0500

"\frac{4-2}{3}\ne \frac{2+1}{4}\ne \frac{3-2}{3}"

Answer: false.